# Fail Fast, Win Big: Rethinking the Drafting Strategy in Speculative Decoding via Diffusion LLMs

### v2

@@ Abstract
Diffusion Large Language Models (dLLMs) offer fast, parallel token generation, but their standalone use is plagued by an inherent efficiency-quality tradeoff. We show that, if carefully applied, the attributes of dLLMs can actually be a strength for drafters in speculative decoding with autoregressive (AR) verifiers. Our core insight is that dLLM’s speed from parallel decoding drastically lowers the risk of costly rejections, providing a practical mechanism to effectively realize the (elusive) lengthy drafts that lead to large speedups with speculative decoding. We present FailFast, a dLLM-based speculative decoding framework that realizes this approach by dynamically adapting its speculation length. It “fails fast” by spending minimal compute in hard-to-speculate regions to shrink speculation latency and “wins big” by aggressively extending draft lengths in easier regions to reduce verification latency (in many cases, speculating and accepting 70 tokens at a time!). Without any fine-tuning, FailFast delivers lossless acceleration of AR LLMs and achieves up to 4.9 speedup over vanilla decoding, 1.7 over the best naive dLLM drafter, and 2.0 over EAGLE-3 across diverse models and workloads. We open-source FailFast at https://github.com/ruipeterpan/failfast.
@@ 1 Introduction
A new wave of Diffusion Large Language Models (dLLMs) (Khanna et al., 2025; Song et al., 2025; Nie et al., 2025; Bie et al., 2025; Ye et al., 2025; Wu et al., 2025b) has emerged as a compelling alternative to the standard autoregressive paradigm in large language models. Unlike autoregressive (AR) LLMs, which are constrained to generating tokens one by one from left to right, dLLMs possess the unique capability to predict and unmask multiple tokens at arbitrary positions simultaneously. Crucially, this decoding process is highly customizable: the model’s unmasking strategy determines exactly which and how many tokens are unmasked during each denoising step (a model forward pass). As such, dLLMs are highly attractive for low-latency inference.
@@
While existing work strives to alleviate the stark compute-accuracy tradeoff of dLLMs as standalone generators (Kang et al., 2025; Bie et al., 2025; aup, 2025), this work instead focuses on motivating and realizing a scenario for which we argue that dLLMs are intrinsically beneficial: as draft models in speculative decoding (Leviathan et al., 2023) with autoregressive target models. Our proposal extends beyond a simple drop-in replacement of dLLMs as drafters in existing speculative decoding strategies to reap their latency benefit – indeed, we show later how this can forego substantial benefits that they bring. Instead, our approach is rooted in two key observations that challenge the status quo for both speculative decoding design and considerations around the limitations of dLLMs.
First, dLLMs can generate long drafts quickly. In AR LLM inference, the decoding latency scales with the number of output tokens (i.e., number of model forward passes) (Yan et al., 2024; Agrawal et al., 2024), whereas dLLMs can unmask multiple tokens in each forward pass, so the latency is instead linear to the number of model forward passes. dLLM’s ability to generate more tokens quickly motivates a rethink of the central challenge involved in designing effective speculative decoding strategies, i.e., balancing that getting longer drafts accepted yields bigger wins (Liu et al., 2024b; Huang et al., 2024), but longer drafts come with higher risk due to each token’s probability of acceptance dropping exponentially. Second, dLLMs exhibit distinct accuracy-compute concavity patterns at the sequence- and draft-level. On a sequence level, the accuracy improvement of each additional forward pass yields diminishing returns. However, within each sequence, there are easier regions of tokens – e.g., simpler tasks like summarizing prior context (Pan et al., 2025a) – where minimal compute (i.e., a single forward pass) suffices for accurate generation. In contrast, for harder regions where existing speculative decoding strategies typically struggle (e.g., difficult tasks like complex arithmetic), dLLMs require more compute to slowly refine their quality. This observation drastically relaxes the issues with the latency-accuracy tension inherent to dLLMs: while additional forward passes are necessary for standalone generation where each token matters for end-to-end quality, they are often unneeded for drafting in speculative decoding. Indeed, beyond the fact that all drafts are ultimately verified (making some errors tolerable), most wins typically come from the easier regions in a sequence (Yan et al., 2024) where initial dLLM draft accuracy routinely suffices.
@@
Across diverse models and workloads, FailFast achieves a speedup of 1.9-4.9 over vanilla decoding, 1.2-1.7 over a strong dLLM drafter baseline, and 1.1-2.0 over EAGLE-3. FailFast achieves lossless acceleration of autoregressive target model generation while employing off-the-shelf dLLMs as draft models with no additional training/fine-tuning.
@@ 2 Background
Diffusion language models. Recent research has established Diffusion Language Models (dLLMs) as a promising alternative to standard autoregressive generation, with prominent examples including both open-source models (Nie et al., 2025; Bie et al., 2025; Ye et al., 2025) and closed-source offerings (Khanna et al., 2025; gem, 2025; Song et al., 2025). Architecturally, dLLMs retain the standard Transformer backbone, preserving access to familiar metrics such as log-probabilities and token confidence scores. However, they diverge fundamentally in their decoding mechanism: rather than generating tokens strictly left-to-right, dLLMs generate text through an iterative unmasking process. This decoding is typically semi-autoregressive (Arriola et al., 2025; gem, 2025; Wu et al., 2025a, b): the output sequence is divided into blocks where inter-block attention remains causal (later blocks attend to earlier blocks), but intra-block attention is bidirectional (this semi-autoregressive decoding also enables optimizations like KV caching (Wu et al., 2025b)). Within each block, tokens are unmasked in a non-deterministic order determined by token confidence rather than position, allowing the model to unmask “easy” tokens before “hard” ones. To improve the efficiency of this iterative process, recent works have proposed acceleration techniques such as approximate KV caching (Wu et al., 2025b; Ma et al., 2025; Liu et al., 2025b), confidence-aware parallel decoding (Wu et al., 2025a, b), self-speculative decoding (Agrawal et al., 2025), etc.
Speculative decoding. Borrowing from classical computer architecture principles (Burton, 1985), speculative decoding has emerged as a standard technique to alleviate the memory-bound nature (Agrawal et al., 2024) of LLM inference (Stern et al., 2018; Leviathan et al., 2023; Yan et al., 2024; Liu et al., 2025a). The process operates on a propose-and-verify cycle: in each round, a lightweight drafter first generates a draft, which is subsequently validated by the target model. The speculation phase prioritizes efficiency, relying on methods such as a standalone draft model (Leviathan et al., 2023; Miao et al., 2024), a trainable module on top of the target model (Cai et al., 2024; Li et al., 2025), a tree-based token cache (Oliaro et al., 2024; Luo et al., 2024; Zhao et al., 2024), an n-gram lookup table (Fu et al., 2024), or a retrieval-based datastore (He et al., 2023). In the verification phase, the target model performs a parallel chunked-prefill over these candidates, which usually consists of either a single sequence of tokens as in (Leviathan et al., 2023) or tree-like structures to further boost the accuracy of speculation (Miao et al., 2024; Cai et al., 2024; Li et al., 2025; Chen et al., 2024), and accepts the longest valid prefix. Consequently, the speculation length is typically conservative (e.g., 3-10) (Liu et al., 2024b; Fu et al., 2024) to maintain an optimal balance between the speculation overhead and the rate of token acceptance.
@@ 3 Motivation
The quality of a dLLM’s output is a function of compute. As non-autoregressive models, dLLMs generate multiple tokens in parallel in an iterative denoising process. In each forward pass, a select few mask tokens with higher confidence within the full sequence of regular and mask tokens are unmasked into regular tokens. As such, the computational cost of a dLLM is determined by the number of denoising steps (forward passes), not the number of tokens generated. By increasing the number of denoising steps, the model iteratively refines the output, trading latency for higher quality/fidelity.
dLLMs can outperform AR LLMs in generation speed while maintaining accuracy by spending fewer forward passes to generate the same number of output tokens. However, naively deploying them as if they are equally-accurate but faster AR drafters in speculative decoding doesn’t capitalize on their full latency benefit. Crucially, we find that dLLMs show concavity in accuracy wins as more compute is spent – improvements with each new forward pass has diminishing returns. In Fig. 2, we naively adopt a dLLM as the drafter and show that a higher compute budget leads to diminishing returns in overall acceptance rate. A dLLM can, in theory, unmask a sequence of unlimited length using a single forward pass. Although the quality of such one-step generation is relatively low and more denoising steps improve the quality – and those denoising steps are crucial for quality if the dLLM is doing standalone generation – the role of draft models in speculative decoding is, by definition, to correctly decode the easier tokens, which is a nice match for the fast yet inaccurate one-step generation.
@@ 3.2 The Varying Decoding Difficulty within Sequences
Importantly, even though the concavity property manifests across entire sequences, within each sequence, there are easier regions and tokens where minimal compute suffices for high-quality generation. Similar to prior work (Liu et al., 2024a; Huang et al., 2024; Pan et al., 2025a), we observe that in natural language generation, the difficulty of generation and speculation differs between regions of the output sequence. Easier regions often consist of syntactic copying, summarization of prior context and input prompt, formulaic enumerations, simple arithmetics, etc. In these regions of lower difficulty, a capable draft model can achieve a near-perfect acceptance rate. Conversely, harder regions involve complex planning, multi-step reasoning, or knowledge retrieval where the draft model’s capacity is insufficient, leading to frequent rejections. In Fig. 3, we pick five random queries from our evaluation datasets, run an AR draft model, and classify tokens in the output sequence into “easier” tokens (including both correct speculations and the “bonus tokens” when all drafted tokens are accepted) and “harder” tokens (rejected by the target model).
@@
Standard speculative decoding approaches typically utilize a static speculation length (e.g., 10 tokens). While this fixed length strives to balance speculation overhead and potential speedup, it is suboptimal for the dynamic nature of generation:
@@
In Easier Regions (e.g., tokens 150-380 for query 4): A short, fixed speculation length results in undue verification overhead. Even though the drafter could correctly speculate the next 10 tokens, its token generation is forced to pause every 10 tokens to invoke the target model for verification. Since target model inference is memory-bound, the frequent loading of model weights incurs significant latency overhead, preventing the system from reaching peak overall efficiency.
@@
In Harder Regions (e.g., tokens 200-250 for query 18): A fixed speculation length results in undue speculation overhead. The drafter wastes compute generating 10 tokens, most of which are likely to be rejected (after the first or second position), resulting in wasted computation and hurting the end-to-end speedup.
@@
Altogether, these properties provide a unique opportunity to change the way drafters are used in speculative decoding. To cope with the varying difficulty of decoding, an ideal speculative decoding paradigm would dynamically reduce the frequency of verification in easy regions to amortize its cost, while minimizing the drafter’s compute in hard regions to minimize wasted effort. Standard autoregressive (AR) draft models are ill-suited for this dynamic strategy – each AR drafter’s forward pass only generates a single token, which makes large speculation lengths (e.g., speculating 30 tokens at a time) prohibitively expensive and risky, since the probability of each token’s acceptance drops exponentially, and the drafting latency eventually overshadows the savings from parallel verification. In contrast, we can reduce the compute budget of dLLM drafters to relax this harsh tradeoff, providing an avenue to get lengthy drafts rapidly.
@@ 4 Method
Capitalizing on the aforementioned observations, we present FailFast, a speculative decoding framework that employs Diffusion LLMs (dLLMs) as draft models to autoregressive verifiers. FailFast departs from traditional optimization strategies by deliberately minimizing the computational effort of the drafter to fail fast and minimize the speculation latency of the draft model; in regions of lower speculation difficulty, it wins big through aggressively extending the speculation lengths and reducing the verification latency of the target model.
@@
There has been a myriad of related work in improving the efficacy of speculative decoding by improving the draft acceptance rate. Typically, this is achieved by enhancing the drafted tokens’ quality through using an ensemble of small drafters to construct token trees (Miao et al., 2024; Cai et al., 2024; Li et al., 2025; Chen et al., 2024), extensive fine-tuning of the draft model (Liu et al., 2023; atl, 2025; Sandler et al., 2025), performing online drafter selection (Liu et al., 2025a), etc.
@@
Easy Regions: In regions of low difficulty (e.g., summarization), even a coarse, 1-step generation from a small dLLM is often sufficient to speculate the correct tokens. Additional refinement steps in these regions are wasted compute.
Hard Regions: In regions of high difficulty (e.g., complex reasoning), a small draft model is statistically likely to diverge from the target model regardless of how many refinement steps are applied. Even though spending extra compute to refine the draft tokens will indeed improve the acceptance rate, it yields diminishing returns, and the quality of speculated tokens is still inherently bounded by the capability of the small drafter, and a significant portion of tokens are eventually rejected anyway.
Thanks to our design choice to adopt dLLMs as draft models, we lower the bar of language modeling quality compared to using dLLMs as standalone models, and can afford to utilize dLLMs in this low-quality yet blazing-fast mode, minimizing the speculation latency per token. We rely on the verification stage to catch errors, ensuring that we “fail fast” and return the task of decoding to the target model with minimal speculation overhead when the task becomes difficult.
@@
While reducing speculation latency addresses the cost of generating drafts, it does not lower the verification latency – the overhead incurred by executing the large target model, which is a factor of both the number of verifications (rounds of speculations) and the verification latency in each round. If we strictly use short speculation lengths, the target model is invoked frequently, creating a latency bottleneck.
@@
To detect these easy regions without ground-truth knowledge of difficulty, we utilize the dLLM’s internal confidence as a proxy of speculation difficulty. Although dLLMs are non-autoregressive, they use the same Transformers backbone as AR LLMs, so we still have access to the probability distribution over the vocabulary for each token. For a dLLM predicting a token at position , we define its confidence as the maximum probability in the output distribution : , where is the vocabulary.
@@
Our adaptive strategy is detailed in Algorithm 1. We begin with a default speculation length . After generating these initial tokens (using very few forward passes), we inspect their confidence scores. If all tokens’ confidence in the current speculated sequence exceeds a confidence threshold (indicating an “easy” region), we assume the draft will likely be accepted. Instead of stopping to verify, we immediately extend the speculation length by another tokens and invoke the drafter again. We repeat this extension process until a low-confidence token is detected or a maximum sequence length is reached. This mechanism allows FailFast to generate massive chunks of tokens (e.g., 60 tokens) in easy regions in a single round of speculation, and getting a majority of them accepted in many cases, which drastically reduces the number of times the target model need to be invoked (verification rounds) – the verification overhead in each round stays the same because short prefills are memory-bound (Agrawal et al., 2024) – while reverting to short bursts in harder regions.
While the notion of dynamically adjusting the draft length is an established concept in speculative decoding (Liu et al., 2024b), the resulting speedups are often limited, hindered by the risk of invoking undue speculation latency. E.g., SpecDec++ (Huang et al., 2024) achieves an average speedup of 10% over baselines. In comparison, FailFast overcomes this by using minimal compute to minimize the risk of incorrect speculations, allowing us to drastically adapt the speculation length on-the-fly.
@@
4.3 Combining FailFast with other optimizations
Our approach is complementary to other optimizations that aim to improve the overall acceptance rate of drafted tokens (Miao et al., 2024; Chen et al., 2024; Liu et al., 2023; atl, 2025; Liu et al., 2025a). Notably, SpecDiff-1 and -2 (Christopher et al., 2025; Sandler et al., 2025) focus on fine-tuning the dLLM drafter for better alignment and a higher acceptance rate, which is complementary to FailFast as the dLLMs we employ can be further fine-tuned to fit the distribution of the base model better. The authors of SpecDiff didn’t release open-source models or their codebase, so we were not able to quantitatively reproduce and compare with them.

### v3

@@ Abstract
Diffusion Large Language Models (dLLMs) offer fast, parallel token generation, but their standalone use is plagued by an inherent efficiency-quality tradeoff. We show that, if carefully applied, the attributes of dLLMs can actually be a strength for drafters in speculative decoding with autoregressive (AR) verifiers. Our core insight is that dLLM’s speed from parallel decoding drastically lowers the risk of costly rejections, providing a practical mechanism to effectively realize the (elusive) lengthy drafts that lead to large speedups with speculative decoding. We present FailFast, a dLLM-based speculative decoding framework that realizes this approach by dynamically adapting its speculation length. It “fails fast” by spending minimal compute in hard-to-speculate regions to shrink speculation latency and “wins big” by aggressively extending draft lengths in easier regions to reduce verification latency (in many cases, speculating and accepting 70 tokens at a time!). Without any fine-tuning, FailFast delivers lossless acceleration of AR LLMs and achieves up to 4.9 speedup over vanilla decoding, 1.7 over the best naive dLLM drafter, and 1.7 over EAGLE-3 across diverse models and workloads. We open-source FailFast at https://github.com/ruipeterpan/failfast.
@@ 1 Introduction
A new wave of Diffusion Large Language Models (dLLMs) (Khanna et al., 2025; Song et al., 2025; gem, 2025; Bie et al., 2025; Wu et al., 2025b) has emerged as a compelling alternative to the standard autoregressive paradigm in large language models. Unlike autoregressive (AR) LLMs, which are constrained to generating tokens one by one from left to right, dLLMs possess the unique capability to predict and unmask multiple tokens at arbitrary positions simultaneously. Crucially, this decoding process is highly customizable: the model’s unmasking strategy determines exactly which and how many tokens are unmasked during each denoising step (a model forward pass). As such, dLLMs are highly attractive for low-latency inference.
@@
While existing work strives to alleviate the stark compute-accuracy tradeoff of dLLMs as standalone generators (Kang et al., 2025; Bie et al., 2025; Qian et al., 2026), this work instead focuses on motivating and realizing a scenario in which we argue that dLLMs are intrinsically beneficial: as draft models in speculative decoding (Leviathan et al., 2023) with autoregressive target models. Our proposal extends beyond a simple drop-in replacement of dLLMs as drafters in existing speculative decoding strategies to reap their latency benefits – indeed, we later show how this can forgo substantial benefits they bring. Instead, our approach is rooted in two key observations that challenge the status quo for both speculative decoding design and considerations around the limitations of dLLMs.
First, dLLMs can generate long drafts quickly. In AR LLM inference, the decoding latency scales with the number of output tokens (i.e., the number of model forward passes) (Yan et al., 2024; Agrawal et al., 2024), whereas dLLMs can unmask multiple tokens in each forward pass, so the latency is instead linear to the number of model forward passes. dLLMs’ ability to generate more tokens quickly motivates a rethink of the central challenge involved in designing effective speculative decoding strategies, i.e., balancing the fact that getting longer drafts accepted yields bigger wins (Liu et al., 2024b; Huang et al., 2024), but longer drafts come with higher risk due to each token’s probability of acceptance dropping exponentially.
Second, dLLMs exhibit distinct accuracy-compute concavity patterns at the sequence- and draft-level. At the sequence level, the accuracy improvement of each additional forward pass yields diminishing returns. However, within each sequence, there are easier regions of tokens – e.g., simpler tasks like summarizing prior context (Pan et al., 2025a) – where minimal compute (i.e., a single forward pass) suffices for accurate generation. In contrast, for harder regions where existing speculative decoding strategies typically struggle (e.g., difficult tasks like complex arithmetic), dLLMs require more compute to slowly refine their quality. This observation drastically relaxes the issues with the latency-accuracy tension inherent to dLLMs: while additional forward passes are necessary for standalone generation, where each token matters for end-to-end quality, they are often unneeded for drafting in speculative decoding. Indeed, beyond the fact that all drafts are ultimately verified (making some errors tolerable), most wins typically come from the easier regions in a sequence (Yan et al., 2024) where initial dLLM draft accuracy routinely suffices.
@@
Across diverse models and workloads, FailFast achieves lossless acceleration of autoregressive target model generation, delivering speedups of up to 4.9 over vanilla decoding, 1.7 over a strong dLLM drafter baseline, and 1.7 over EAGLE-3. We evaluate on diverse models and datasets and show that the intrinsic properties of dLLMs motivate fundamental changes to drafting strategies, which FailFast exploits to leverage their strengths as draft models. By enabling off-the-shelf, already-aligned (Wu et al., 2025a; Bie et al., 2025; Gong et al., 2024; Zhou et al., 2025) dLLMs to act as effective drafters without any additional training or fine-tuning, FailFast fundamentally changes the drafting paradigm and naturally benefits from continued improvements in dLLM alignment (Sandler et al., 2025).
@@ 2 Background
Diffusion language models. Recent research has established Diffusion Language Models (dLLMs) as a promising alternative to standard autoregressive generation, with examples spanning both open-source (Nie et al., 2025; Bie et al., 2025; Ye et al., 2025) and closed-source models (Khanna et al., 2025; gem, 2025; Song et al., 2025). Architecturally, dLLMs retain the standard Transformer backbone, preserving access to familiar metrics such as log-probabilities and token confidence scores. However, they diverge fundamentally in their decoding mechanism: instead of generating tokens strictly left-to-right, dLLMs generate text through an iterative unmasking process. This decoding is typically semi-autoregressive (Arriola et al., 2025; gem, 2025; Wu et al., 2025a, b): the output sequence is divided into blocks, with inter-block attention remaining causal (later blocks attend to earlier blocks), but intra-block attention is bidirectional. (This semi-autoregressive decoding also enables optimizations like KV caching (Wu et al., 2025b).) Within each block, tokens are unmasked in a non-deterministic order determined by token confidence rather than position, allowing the model to unmask “easy” tokens before “hard” ones. To further improve efficiency, recent works have proposed acceleration techniques such as approximate KV caching (Wu et al., 2025b; Ma et al., 2025; Liu et al., 2025b), confidence-aware parallel decoding (Wu et al., 2025a, b), and self-speculative decoding (Agrawal et al., 2025).
Speculative decoding. Borrowing from classical computer architecture principles (Burton, 1985), speculative decoding has emerged as a standard technique to alleviate the memory-bound nature (Agrawal et al., 2024) of LLM inference (Stern et al., 2018; Leviathan et al., 2023; Yan et al., 2024; Liu et al., 2025a). The process operates on a propose-and-verify cycle: in each round, a lightweight drafter first generates a draft, which is subsequently validated by the target model. The speculation phase prioritizes efficiency, relying on methods such as a standalone draft model (Leviathan et al., 2023; Miao et al., 2024), a trainable module on top of the target model (Cai et al., 2024; Li et al., 2025), a tree-based token cache (Oliaro et al., 2024; Luo et al., 2024; Zhao et al., 2024), an n-gram lookup table (Fu et al., 2024), or a retrieval-based datastore (He et al., 2023). In the verification phase, the target model performs a parallel chunked-prefill over these candidates, which usually consists of either a single sequence of tokens as in (Leviathan et al., 2023) or tree-like structures to further boost the quality of speculation (Miao et al., 2024; Cai et al., 2024; Li et al., 2025; Chen et al., 2024), and accepts the longest valid prefix. Consequently, the speculation length is typically conservative (e.g., 3-10) (Liu et al., 2024b; Fu et al., 2024) to maintain an optimal balance between the speculation overhead and the rate of token acceptance.
@@ 3 Motivation
The quality of a dLLM’s output is a function of compute. As non-autoregressive models, dLLMs generate multiple tokens in parallel during an iterative denoising process. In each forward pass, a select few masked tokens with higher confidence within the full sequence of regular and mask tokens are unmasked into regular tokens. As such, the computational cost of a dLLM is predominantly determined by the number of denoising steps (forward passes), not the number of tokens generated. By increasing the number of denoising steps, the model iteratively refines the output, trading latency for higher quality/fidelity.
dLLMs can outperform AR LLMs in generation speed while maintaining accuracy by spending fewer forward passes to generate the same number of output tokens (Wu et al., 2025b). However, naively deploying them as if they are equally-accurate but faster AR drafters in speculative decoding does not capitalize on their full latency benefit. Crucially, we find that dLLMs show concavity in accuracy wins as more compute is spent – improvements with each new forward pass have diminishing returns. In Fig. 2, we naively adopt a dLLM as the drafter and show that a higher compute budget leads to diminishing returns in overall acceptance rate. A dLLM can, in theory, unmask a sequence of unlimited length using a single forward pass. Although the quality of such one-step generation is relatively low and more denoising steps improve the quality – and those denoising steps are crucial for quality if the dLLM is doing standalone generation – the role of draft models in speculative decoding is, by definition, to correctly decode the easier tokens, which is a nice match for the fast yet inaccurate one-step generation.
In Fig. 2, we analyze the theoretical speedup (Theorem 3.8 in (Leviathan et al., 2023)) of speculative decoding over vanilla decoding using different drafters of the same size. The autoregressive drafter generates one token in each forward pass, incurring drafting latency linear in the draft length, whereas the semi-autoregressive dLLM drafter performs one-step generation and generates 8 tokens in each forward pass (Wu et al., 2025a), resulting in sublinear drafting latency. Even though longer speculation lengths exponentially reduce the acceptance probabilities of the tokens later in the draft, the dLLM’s one-step generation compensates by dramatically lowering the average latency to generate each draft token, increasing the maximum speedup from 3.0 (at speculation length 8) to 5.2 (at length 16).
@@ 3.2 The Varying Decoding Difficulty within Sequences
Importantly, even though the concavity property manifests across entire sequences, within each sequence, there are easier regions and tokens where minimal compute suffices for high-quality generation. Similar to prior work (Liu et al., 2024a; Huang et al., 2024; Pan et al., 2025a), we observe that in natural language generation, the difficulty of generation and speculation differs between regions of the output sequence. Easier regions often consist of syntactic copying, summarization of prior context and input prompt, formulaic enumerations, simple arithmetic, etc. In these regions of lower difficulty, a capable draft model can achieve a near-perfect acceptance rate. Conversely, harder regions involve complex planning, multi-step reasoning, or knowledge retrieval where the draft model’s capacity is insufficient, leading to frequent rejections. In Fig. 4, we pick five random queries from one of our evaluation datasets, run an AR draft model, and classify tokens in the output sequence into “easier” tokens (including both correct speculations and the “bonus tokens” when all drafted tokens are accepted) and “harder” tokens (rejected by the target model). Further, in Fig. 5, we report the fraction of tokens that belong to a consecutive block of more than easy tokens across different tasks. If an easy region is defined to be a consecutive block of 10 “easy” tokens, 50-70% of tokens belong to an easy region. Moreover, 15-26% of tokens belong to an easy region of 50 tokens.
@@
Standard speculative decoding approaches typically utilize a static speculation length (e.g., tokens). While this aims to balance speculation overhead and potential speedup, it is suboptimal for the dynamic nature of generation:
@@
In easier regions (e.g., tokens 150-380 for query 4 in Fig. 4): A short, fixed speculation length results in undue verification overhead. Even though the drafter could correctly speculate the next 10 tokens, its token generation is forced to pause every 10 tokens to invoke the target model for verification. Since target model inference is memory-bound, the frequent loading of model weights incurs significant latency overhead, preventing the system from reaching peak overall efficiency.
@@
In harder regions (e.g., tokens 200-250 for query 18 in Fig. 4): A fixed speculation length results in undue speculation overhead. The drafter wastes compute generating 10 tokens, most of which are likely to be rejected (after the first or second position), resulting in wasted computation and hurting the end-to-end speedup.
@@
Altogether, these properties provide a unique opportunity to change the way drafters are used in speculative decoding. To cope with the varying difficulty of decoding, an ideal speculative decoding paradigm would dynamically increase the speculation length in easy regions to reduce the frequency of verification and amortize its cost, while minimizing the drafter’s compute in hard regions to minimize wasted effort. Standard autoregressive (AR) draft models are ill-suited for this dynamic strategy – each AR drafter’s forward pass only generates a single token, which makes large speculation lengths (e.g., speculating 30 tokens at a time) prohibitively expensive and risky (Fig. 2), since the probability of each token’s acceptance drops exponentially, and the drafting latency eventually overshadows the savings from parallel verification. In contrast, we can reduce the compute budget of dLLM drafters to relax this harsh tradeoff, providing an avenue to get lengthy drafts rapidly.
@@ 4 Method
Capitalizing on the aforementioned observations, we present FailFast, a speculative decoding framework that employs Diffusion LLMs (dLLMs) as draft models to autoregressive verifiers. FailFast departs from traditional optimization strategies by deliberately minimizing the computational effort of the drafter to fail fast and minimize the speculation latency of the draft model; in regions of lower speculation difficulty, it wins big through aggressively extending the speculation length and reducing the verification latency of the target model.
@@
There has been a myriad of related work on improving the efficacy of speculative decoding by improving the draft acceptance rate. Typically, this is achieved by enhancing the drafted tokens’ quality through using an ensemble of small drafters to construct token trees (Miao et al., 2024; Cai et al., 2024; Li et al., 2025; Chen et al., 2024), extensive fine-tuning of the draft model (Liu et al., 2023; atl, 2025; Sandler et al., 2025), performing online drafter selection (Liu et al., 2025a), etc.
@@
Easy regions: In regions of low difficulty (e.g., summarization), even a coarse, one-step generation from a small dLLM is often sufficient to speculate the correct tokens (Fig. 2). Additional refinement steps in these regions are wasted compute.
Hard regions: In regions of high difficulty (e.g., complex reasoning), a small draft model is statistically likely to diverge from the target model regardless of how many refinement steps are applied. Although spending additional compute to refine the draft tokens does indeed improve the acceptance rate, it yields diminishing returns (Fig. 2): the quality of speculated tokens is still inherently bounded by the capability of the small drafter, and a significant portion of tokens is eventually rejected anyway.
By adopting dLLMs as draft models, we relax the requirement for high-quality language modeling compared to using them as standalone models, allowing us to operate in a low-quality but blazing-fast mode that minimizes speculation latency per token. The verification stage then catches errors, enabling a “fail fast” approach that quickly returns the task of decoding to the target model with minimal speculation overhead when the task becomes difficult.
@@
While reducing speculation latency addresses the cost of generating drafts, it does not lower the verification latency – the overhead incurred by executing the large target model, which is a factor of both the number of verifications (rounds of speculation) and the verification latency in each round. If we strictly use short speculation lengths, the target model is invoked frequently, creating a latency bottleneck.
@@
To detect these easy regions without ground-truth knowledge of difficulty, we utilize the dLLM’s internal confidence (Fu et al., 2025; Dai et al., 2024) as a proxy for speculation difficulty. Although dLLMs are non-autoregressive, they use the same Transformers backbone as AR LLMs, so we still have access to the probability distribution over the vocabulary for each token. For a dLLM predicting a token at position , we define its confidence as the maximum probability in the output distribution : , where is the vocabulary.
@@
Our adaptive strategy is detailed in Algorithm 1. We begin with a default speculation length . After generating these initial tokens (using minimal compute), we inspect their confidence scores. If all tokens’ confidence in the current speculated sequence exceeds a confidence threshold (indicating an “easy” region), we assume the draft will likely be accepted. Instead of stopping to verify, we immediately extend the speculation length by another tokens and invoke the drafter again. We repeat this extension process until a low-confidence token is detected or a maximum sequence length is reached. This mechanism allows FailFast to generate massive chunks of tokens (e.g., 70 tokens) in easy regions in a single round of speculation, and to get a majority of them accepted in many cases, which drastically reduces the number of times the target model needs to be invoked (verification rounds) – the verification overhead in each round stays the same because short prefills are memory-bound (Agrawal et al., 2024) – while reverting to short bursts in harder regions.
While the notion of dynamically adjusting the draft length is an established concept in speculative decoding (Liu et al., 2024b), the resulting speedups are often limited, hindered by the risk of invoking undue speculation latency. For example, SpecDec++ (Huang et al., 2024) achieves an average speedup of 10% over its baselines. In comparison, FailFast overcomes this by using minimal compute to minimize the risk of incorrect speculation, allowing it to drastically adapt the speculation length on-the-fly.
@@
4.3 Policy Design vs. Training-Based Approaches
Our work focuses specifically on how to adapt drafting policies when using dLLMs as drafters, rather than improving performance through additional training. Indeed, most state-of-the-art dLLMs are adapted from pretrained autoregressive (AR) weights (Wu et al., 2025a; Bie et al., 2025; Gong et al., 2024; Zhou et al., 2025) and are therefore already implicitly aligned with the target AR models when used as drafters. Leveraging this property, we study how to better exploit dLLMs at inference time via policy design. That said, our approach is complementary to prior optimizations that aim to improve the overall acceptance rate of drafted tokens (Miao et al., 2024; Chen et al., 2024; Liu et al., 2023; atl, 2025; Liu et al., 2025a), including recent work that fine-tunes the dLLM drafter for improved alignment and higher acceptance rates (Christopher et al., 2025; Sandler et al., 2025). These techniques are complementary to FailFast, as the dLLMs we employ can be further fine-tuned to fit the base model’s distribution better.